# Learning Spatially Adaptive $\ell_{1}$-Norms Weights for Convolutional Synthesis Regularization

### v3

@@ Abstract
Index Terms:
Neural Networks, Convolutional Dictionary Learning, Sparsity, Adaptive Regularization, Low-Field MRI
@@ I Introduction
$\frac{\alpha}{2}$ $\sum_{k=1}^{K}\|\mathbf{x}-d_{k}\ast\mathbf{s}_{k}\|_{2}^{2}+\lambda\sum_{k=1}^{K}\|\mathbf{s}_{k}\|_{1}$
@@ III-C Training and Implementation Details
The 2D U-Net $\mathbf{u}_{\Theta}$ used in the CNN-block $\mathrm{NET}_{\Theta}$ in (9) has a 2-$K$-$2K$-$4K$-$2K$-$K$ structure with two input-channels (for the real and imaginary parts of the image) and $K$ output channels for $K$ parameter maps $\left\{\mathbf{\Lambda}_{k}\right\}_{k}$. To handle complex-valued data, note that the real and the imaginary parts of the images are approximated using the same filters, i.e. we have $\mathbf{D}\mathbf{s}:=\sum_{k}(d_{k}\ast\mathrm{Re}(\mathbf{s}_{k})+\mathrm{i}\,d_{k}\ast\mathrm{Im}(\mathbf{s}_{k}))$. For training, we used the Adam optimizer [20] with an MSE-loss using a batch-size of 2. For all methods, the scalar regularization parameters ($\lambda$ and $\beta$), which were ”activated” using a Softplus activation function to maintain strict positivity, were assigned a learning rate of $10^{-1}$, while the U-Nets $\mathbf{u}_{\Theta}$ were given a learning rate of $10^{-4}$ with a weight decay regularization weight of $10^{-5}$.
The dictionary filters $\{d_{k}\}_{k}$ were pre-trained on 160 randomly selected images from the training set using the algorithm described in [21] implemented in SPORCO [22]. We tested different kernel sizes $k_{f}\times k_{f}=7\times 7$, $11\times 1$ 11 and $13\times 13$ as well as different numbers of filters $K=32,64$. Due to space limitations, we here only report results for $K=64$ and $k_{f}\times k_{f}=11\times 11$. We point out that the experiments suggest that more and larger filters tend to improve the achievable performance. The upper bound $t>0$ in (9) was fixed to $t=10.0$, which is also the average maximal magnitude value the images were normalized to.

### v4

@@ I Introduction
$\frac{\alpha}{2}$ $\|\mathbf{x}-\sum_{k=1}^{K}d_{k}\ast\mathbf{s}_{k}\|_{2}^{2}+\lambda\sum_{k=1}^{K}\|\mathbf{s}_{k}\|_{1}$
@@ III-C Training and Implementation Details
The 2D U-Net $\mathbf{u}_{\Theta}$ used in the CNN-block $\mathrm{NET}_{\Theta}$ in (9) has a 2-KK-2K2K-4K4K-$2K$-$K$ structure with two input-channels (for the real and imaginary parts of the image) and $K$ output channels for $K$ parameter maps $\left\{\mathbf{\Lambda}_{k}\right\}_{k}$. To handle complex-valued data, note that the real and the imaginary parts of the images are approximated using the same filters, i.e. we have $\mathbf{D}\mathbf{s}:=\sum_{k}(d_{k}\ast\mathrm{Re}(\mathbf{s}_{k})+\mathrm{i}\,d_{k}\ast\mathrm{Im}(\mathbf{s}_{k}))$. For training, we used the Adam optimizer [20] with an MSE-loss using a batch-size of 2. For all methods, the scalar regularization parameters ($\lambda$ and $\beta$), which were ”activated” using a Softplus activation function to maintain strict positivity, were assigned a learning rate of $10^{-1}$, while the U-Nets $\mathbf{u}_{\Theta}$ were given a learning rate of $10^{-4}$ with a weight decay regularization weight of $10^{-5}$.
The dictionary filters $\{d_{k}\}_{k}$ were pre-trained on 160 randomly selected images from the training set using the algorithm described in [21] implemented in SPORCO [22]. We tested different kernel sizes $k_{f}\times k_{f}=7\times 7$, $11\times 11$ and $13\times 13$ as well as different numbers of filters $K=32,64$. Due to space limitations, we here only report results for $K=64$ and $k_{f}\times k_{f}=11\times 11$. We point out that the experiments suggest that more and larger filters tend to improve the achievable performance. The upper bound $t>0$ in (9) was fixed to $t=10.0$, which is also the average maximal magnitude value the images were normalized to.